OpenStudy (anonymous):

If a salesman earned a 6% commission on sales, and in one week he earned $300 in commissions, how much merchandise did he sell that week?

OpenStudy (lucaz):

6% = 6/100 = 0.06 6%(x)=$300 x=300/0.06 x=$5,000

OpenStudy (lucaz):

I think it's like this

OpenStudy (anonymous):

Thanks Lucaz!

OpenStudy (lucaz):

=]